OpenStudy (anonymous):

What is: Area of a circle with 10cm diameter? Perimeter of a semicircle with radius 14cm? The radius of a circle whose circumference is 80cm?

OpenStudy (anonymous):

area of i circle: 25*Pi perimeter: (2*Pi*14)/2=14Pi radius: 2Pi*r=80 r=80/2Pi r=40Pi

OpenStudy (anonymous):

Area of a circle is given by \(\pi r^{2}\), or \(\frac{1}{2}\tau \: r^{2}\) = \(78.5 cm^{2}\) Perimeter of a circle is given by \(2 \pi r\) or \(\tau \: r\) = \(44.0 cm\) Radius is given by \(\frac{c}{2\pi}\) or \(\frac{c}{\tau}\) = \(12.7cm\)